A car is moving at 65.0 km/h at a distance of 125 m from a traffic light when the light turns red. The weight of the car is 14800 N


(a) What is the mass of the car in kg?


(b) If the driver wishes to stop at the light, what is the required constant acceleration of the car in m/s^2


(c) What will be the velocity of the stone when it hits the water?




1
Expert's answer
2021-02-10T12:20:57-0500

a)


"m=\frac{W}{g}=\frac{14800}{9.8}=1510\ kg"

b)


"v^2=2as\\\frac{65^2}{3.6^2}=2(125)a\\a=1.30\frac{m}{s^2}"

c)


"F_f=ma=(1510)(1.3)=1960\ N"